OpenStudy (anonymous):

Soybean meal is 14% protein; cornmeal is 7% protein. How many pounds of each should be mixed together in order to get 280-lb mixture that is 9% protein?

OpenStudy (amistre64):

we could work it all out or just use a compact formula; ill use the formula :) soy = 280(.09-.07)/(.14-.07)

OpenStudy (amistre64):

and corn = 280 - soy

OpenStudy (amistre64):

i get 80 soy leaving 100 for corn; but you might wanna dbl chk

OpenStudy (amistre64):

..200 for corn lol

OpenStudy (lalaly):

s+c=280 The total amount of protein must equal 9% of 280, 0.14s + 0.07c = 0.09(280) solve using elimination or substitution

OpenStudy (amistre64):

we dbl chk by multiplying the amount(%)s and see if they match 80(.14) + 200(.07) = 280(.09) 11.2 + 14 = 25.2 25.2 = 25.2

OpenStudy (anonymous):

so it would be 80 lb for soybean & 100 lb for Cornmeal?

OpenStudy (amistre64):

80 for soy; which leaves 200 for corn ... for a total of 280

OpenStudy (anonymous):

ok